Calculate the temperature at which root mean square speed of a gas molecules is double of its speed at 27°C pressure remaining constant

1
Expert's answer
2022-02-15T16:19:55-0500

The rms speed (u) is directly proportional to the square root of temperature.

"\frac{u'}{u}=\sqrt{\frac{T'}{T}}"

"\frac{2u}{u}=\sqrt{\frac{T'}{27+273}}"

"2=\sqrt{\frac{T'}{300}}"

"4=\frac{T'}{300}"

"T'=1200K"

Thus, at 1200 K would the rms speed of a gas molecule have twice its value at 27 deg C